OpenStudy (anonymous):

Solve for x. 2 • 3 – 2 • 11 = 2(x – 11) A. 3 B. 9 C. 6 D. 2

OpenStudy (the_fizicx99):

2 • 3 – 2 • 11 = 2(x – 11) 2 * 3 - 2 * 11 = 2x -22 Can you solve now?

OpenStudy (the_fizicx99):

*cough* reported

OpenStudy (the_fizicx99):

@freakenpurplepizza can you solve? :o

OpenStudy (anonymous):

3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

2 * 3 - 2 * 11 = 2x -22 6 - 22 = 2x - 22 +22 +22 6 = 2x /2 /2 Yesh its 3 :)

OpenStudy (the_fizicx99):

medal ;~~~;

OpenStudy (anonymous):

there you go!! @tHe_FiZiCx99

OpenStudy (the_fizicx99):

yishh :)